OpenStudy (anonymous):

Fiber-optic cables are used widely for internet wiring, data transmission, and surgeries. When light passes through a fiber-optic cable, its intensity decreases with the increase in the length of the cable. If 1500 lumens of light enters the cable, the intensity of light decreases by 3.4% per meter of the cable. Can this situation be represented by a linear function? Write a function f(x) to represent the intensity of light, in lumens, when it has passed through x meters of the cable.

OpenStudy (anonymous):

Some scientists are trying to make a cable for which the intensity of light would decrease by 5 lumens per unit length of the cable. Can this situation be represented by a linear function? Justify your answer and write the appropriate function to represent this situation if 1500 lumens of light enter the cable

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (kropot72):

a) The intensity of light after passing through 1 meter of cable is given by: \[\large I _{1}=1500\times(1.000-0.034)=1500\times0.966\] What will the intensity be after passing through 2 meters of cable? \[\large I _{2}=?\]

OpenStudy (anonymous):

1500*(1-0.034)^2

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

You are correct. And what will the intensity be after passing through x meters of cable?

OpenStudy (anonymous):

I don't know, how would i do that?

OpenStudy (anonymous):

can it be represented by a linear function or not?

OpenStudy (kropot72):

\[\large I _{1}=1500\times0.966\] \[\large I _{2}=1500\times0.966^{2}\] \[\large I _{3}=1500\times0.966^{3}\] \[\large I _{4}=1500\times0.966^{4}\] \[\large I _{x}=1500\times......?\]

OpenStudy (anonymous):

oh, I got it, the intensity will be increasing at the same value each meter, which can be explained by finding the rate of change.

OpenStudy (kropot72):

Not really. Calculating the intensity after 1, 2 and 3 meters gives: \[\large I _{1}=1449\] \[\large I _{2}=1399.734\] \[\large I _{3}=1352.143\] \[\large I _{1}-I _{2}=49.266\] \[\large I _{2}-I _{3}=47.591\] The required equation will not be linear.

OpenStudy (kropot72):

The required equation can be seen by looking at the pattern that I posted previously. It is as follows: \[\large f(x)=1500\times0.966^{x}\]

OpenStudy (anonymous):

what is the .966 from??

OpenStudy (kropot72):

b) In the second case the intensity decreases by a fixed amount per unit length. Therefore the equation will be linear, the reason being there will be no higher powers of the variable higher than unity.

OpenStudy (anonymous):

hmm but wouldn't it be 1.034, because f(x)=p(1+r)^x

OpenStudy (anonymous):

but if its decreasing I guess it would be .966

OpenStudy (kropot72):

"what is the .966 from??" The intensity in part a) decreases by 3.4% per meter. Therefore the intensity after 1 meter will be: \[\large I _{1}=1500\times(1.000-0.034)=1500\times0.966\]

OpenStudy (anonymous):

k gotcha, so the function is f(x)=1500(1-.034)^x

OpenStudy (kropot72):

Yes, you are correct for part a) The equation for part b) the equation is: \[\large f(x)=1500-5x\]

OpenStudy (anonymous):

So that would be for the scientist one?

OpenStudy (anonymous):

And that is linear

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

thanks!!